# Diffraction Aided Wireless Positioning

### v1

@@ IV-A TOF range measurements
Here, $\gamma_{k}(\theta_{tx,k},\theta_{rx,k})$ is the complex gain for the received signal due to the $k^{th}$ MPC and includes the effects of the transmit and receive side antennas, transmitted signal power and also the path gain $h_{k}$. For further analysis, the receiver noise $n(t)$ is assumed to be Gaussian. Out of the $K$ MPCs in the received signal, let indices $k=0$ and $k=1$ correspond to the two diffraction MPCs. Therefore, the remaining $K$-$2$ MPCs consist of those MPCs that contain more than one interaction with the environment. If the electric field for the transmit signal is parallel to the horizontal diffracting window edges, Theorem 1 applies. Therefore, it leads to the upper edge diffraction MPC dominating over the lower edge diffraction MPC. In a system this could be achieved by mounting the linear polarized antennas horizontally, instead of the conventional vertically mounted implementation. Thus, the received signal can be written as
$\begin{split}&r(t)\approx\underbrace{\gamma_{0}(\theta_{tx,0},\theta_{rx,0})p(t-\tau_{0})}_{dominating\;\;diffraction\;\;MPC}\\ &\hskip 50.00008pt+\underbrace{\sum_{k=2}^{K-1}\gamma_{k}(\theta_{tx,k},\theta_{rx,k})p(t-\tau_{k})}_{MPC-4}+n(t)\\ \end{split}$ (17)
Next, we assume resolvability between the MPCs i.e., $|\tau_{i}-\tau_{j}|>T_{p},\;\;\forall\;\;i\neq j$. It can be shown with this assumption, that the CRLB for estimating the propagation delay of the first arriving path $\tau_{0}$ is given by [35, 36]
$\begin{split}\text{CRLB}(\tau_{0})=\frac{1}{8\pi^{2}\beta^{2}\text{SNR}}=\frac{N_{0}}{8\pi^{2}\beta^{2}(\gamma_{0}(\theta_{rx,0},\theta_{tx,0}))^{2}}.\end{split}$ (18)
Here, $\gamma_{0}(\theta_{rx,0},\theta_{tx,0})$ is the amplitude corresponding to the first arriving path, $SNR$ is receive side signal to noise ratio (linear), and $\beta$ is the bandwidth of the transmitted signal (Hz). Since this first arriving path corresponds to the dominating diffraction path, we have $\tau_{0}=\frac{p}{c}$. Here, $\tau_{0}$ is the propagation delay of the first arriving path, and the path length of the dominating diffraction path $p$ is given by eq. (12) and $c$ is the speed of light. Drawing insights from the raytracing analysis in Appendix -D and eq. (18), we can improve the estimate of $p$ by maximizing the antenna gain for $AoA=90$°. In other words, we could improve the ranging performance if we orient the receiver side antenna such that the antenna gain is maximum towards the windows (or in the horizontal direction). Further improvement to the ranging estimate can be made by scanning the transmit side beam to maximize the receiver side power of the dominant diffracting MPC.
@@ IV-B Fisher Information analysis of the 3D position estimate
In this section, we examine the conditions under which we can perform 3D NLOS localization using the TOF-based ranging estimates corresponding to the dominating diffraction MPC using Fisher Information analysis and then calculate the lower bound to the 3D position error. In Fig. 3, we have ‘$M$’ UAV anchors in front of the building. The node is located on an arbitary building floor which has two diffracting edges separated by a vertical distance $w$. We assume the origin is located in the bottom middle of the front wall of the building at the intersection of the front wall with the the ground the x-axis oriented along the diffracting edges and the Z-axis pointed along the vertical direction as in Fig. 3.
@@
Each anchor indexed by ‘$j$’ is positioned at known coordinates $\bm{X}^{a}_{j}=[x_{a,j},y_{a,j},z_{a,j}]^{T}$ relative to the origin, while the node is located on an arbitrary floor of the building at $\bm{\alpha}=[x_{n},y_{n},z_{n}]^{T}$. Assuming the anchors are transmitting x-polarized signals, each TOF-based ranging measurement corresponds to the path length of the respective dominating diffraction MPC between each anchor $j$ and the node. With a slight change in notation, we denote the path length of the respective dominating diffraction MPC as $p_{j}(\bm{\alpha})$ obtained using eq. (12), with the subscript ‘j’ representing the $j^{th}$ anchor. Consequently, we obtain $M$ noisy ranging measurements from which we need to estimate the 3D position of the node. Assuming the 3D position of the node $\bm{\alpha}$ as the vector of the three parameters to be estimated, its corresponding Fisher Information Matrix (FIM) $\bm{J}_{\bm{\alpha}}\in\mathbb{R}^{3\times 3}$ [37], [38] and can be expressed as
$\bm{J}_{\bm{\alpha}}=c^{2}\bm{J}\bm{J}_{\tau_{0}}\bm{J}^{T}.$ (19)
Since we assume the noise in our range measurements is independent across the $M$ anchors, $\bm{J}_{\tau_{0}}=diag(\lambda_{0},\cdots,\lambda_{M-1})$ is a diagonal matrix where $\lambda_{j}=\frac{1}{8\pi^{2}\beta^{2}SNR_{j}}$ represents the inverse of the CRLB of the path delay estimate of the first arriving path between the $j^{th}$ anchor and the node. The Jacobian $\bm{J}\in\mathbb{R}^{3\times M}$ is the transform from the ranging measurement space $\in\mathbb{R}^{M}$ to the unknown parameter to be estimated space $\in\mathbb{R}^{3}$ and is given by
$\bm{J}=\begin{bmatrix}\frac{\partial{p_{0}(\bm{\alpha})}}{\partial x_{n}}&\cdots&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial x_{n}}\\ \frac{\partial{p_{0}(\bm{\alpha})}}{\partial y_{n}}&\cdots&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial y_{n}}\\ \frac{\partial{p_{0}(\bm{\alpha})}}{\partial z_{n}}&\cdots&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial z_{n}}\end{bmatrix}.$ (20)
@@
The FIM $\bm{J}_{\bm{\alpha}}\in\mathbb{R}^{3\times 3}$ defined in eq. (19) is invertible if and only if $rank(\bm{J})=3$. Here, the Jacobian Matrix $\bm{J}\in\mathbb{R}^{3\times M}$ is defined in eq. (20) from the TOF ranging measurement space $\mathbb{R}^{M}$ to the 3D coordinate space $\bm{\alpha}\in\mathbb{R}^{3}$.
@@
$\sqrt{\bm{A}}$ is the elementwise square root operation of a matrix $\bm{A}$. We also know that multiplying a matrix with a non-zero diagonal matrix does not change its rank. Further, by using $rank(\bm{A}\bm{A}^{T})=rank(\bm{A})$ [40], we obtain $rank(\bm{J}_{\bm{\alpha}})=rank(\bm{J})$. Now, $\bm{J}_{\bm{\alpha}}$ is invertible if and only if $rank(\bm{J}_{\bm{\alpha}})=3$. Hence, we obtain the desired result. ∎
@@
  • It is necessary but not sufficient to have at least three anchors. This is because, for $rank(\bm{J})=3$, it is necessary but not sufficient to have $M\geq 3$.
@@
To illustrate that having $M=3$ anchors is not a sufficient condition for 3D positioning, consider a scenario where we have three anchors such that their x-coordinate $x_{a,j}=0\;\;\forall\;\;j\in[0,2]$. Observe that as the node position is varied inside the building such that it approaches the plane $X=0$, i.e., $x_{n}\to 0$, for each anchor, we have the x-coordinate of the corresponding diffraction point $q_{x}\to 0$. Notice in the first row of the Jacobian $J$, defined in eq. (-E), the term $\frac{\partial p_{j}{(\bm{\alpha}})}{\partial x_{n}}\to 0\;\;\forall j\in[0,2]$. Therefore, for $x_{n}=0$, $rank(J)\neq 3$, and by using Theorem 2, we cannot estimate the 3D position of the node for this scenario.
The diffracting edges are located parallel to the X-axis at $Y=0$. Observe the x-coordinates of the end points of the diffracting edges correspond to the width of the building which is required to be known a priori whereas the height of the diffracting edges is not required to be known. This is because the node and the edge from which diffraction takes place are located on the same building floor. Therefore, by estimating the node location leads to an estimate of the height of the edge since the diffracting edge is located at a distance $\frac{w}{2}$ above the node. Note corollary 3 leads to the following observation - we need position knowledge of the anchors in local building coordinates. These local coordinates are defined with respect to the origin located on the diffracting edges as shown in Fig. 3.
@@
V Numerical Results
In this section, we lower bound the achievable performance by numerically simulating the analytical CRLB expression in previous sections for a public safety scenario. We demonstrate both the 3D and Z-axis positioning performance in NLOS scenarios can be improved with our positioning technique. A Linear Least Squares (LLS) technique [38] estimation technique is used as a baseline comparison to our proposed positioning technique. Previously, LLS has been shown to achieve its CRLB under the assumption that the Gaussian noise in ranging measurements is much smaller than the path lengths between the anchors and the node [41]. The baseline comparison is a model mismatch scenario, where for the NLOS scenario, the ranging measurements are noisy measurements of the the diffraction paths, whereas, LLS assumes Euclidean path lengths.
For the simulation, each building floor contains a window and we consider $M=3$ anchors located on one side of the building at fixed known coordinates. The numerical values for the other simulation parameters are shown in Table V. As a performance metric, we look at the Position Error Bound (PEB) [39], which is a lower bound on the position estimate of any unbiased estimator. The 3D PEB is defined to be $\sqrt{\mathbb{T}\{(\bm{J}^{-1}_{\bm{\alpha}})\}}$ where the operator ${.}\mathbb{T}\{.\}$ is the trace of a square matrix and $\bm{J}^{-1}_{\bm{\alpha}}$ is the inverse of the FIM defined in eq. (19). Similarly, the bound on the Z-axis is defined to be $\sqrt{{(\bm{J}^{-1}_{\bm{\alpha}})}^{[3,3]}}$, where the notation $\bm{A}^{[i,j]}$ represents the element in the $i^{th}$ row and $j^{th}$ column of matrix $\bm{A}$.
Finally, we uniformly sample $100,000$ points representing candidate node locations $\bm{\alpha}$ within a cuboidal shaped building. For the simulation we consider a cuboidal volume such that $[-10,0,5]^{T}\leq\bm{\alpha}\leq[10,20,40]^{T}$, where the $\leq$ operator between two vectors represents an element-wise operation. Thus, we obtain the 3D-PEB and Z-PEB values for possible node locations within the building. The final results are plotted as a cumulative density histogram in Fig. 4(a) and Fig. 4(b) respectively. The baseline positioning method despite the high SNR is unable to deal with NLOS bias in the ranging measurements and thus offers poor performance. Positioning based on ranging measurements using the diffraction path model hugely improves both the 3D and Z-axis positioning estimates by directly estimating the NLOS path length and is only limited in performance due to system parameters like signal bandwidth and noise figure.
@@ V Numerical Results
Anchor 00 pos. $\bm{X}_{0}^{a}=[-10m,-20m,-10m]^{T}$
Anchor $1$ pos. $\bm{X}_{1}^{a}=[0m,-7m,-20m]^{T}$
Anchor $2$ pos. $\bm{X}_{2}^{a}=[10m,-20m,-10m]^{T}$
Building endpoint $1$ $x_{1}=-10m$
Building endpoint $2$ $x_{2}=10m$
window height $w=2m$
Signal Bandwidth $\beta=100$ MHz
Signal-to-Noise Ratio SNR $\in[3dB,6dB,9dB,12dB,15dB]$
(a) 3D Positioning Performance
(b) $Z$-axis Positioning Performance
Figure 4: Positioning Performance comparison of the proposed diffraction inspired positioning technique vs the baseline. The NLOS diffraction path model mitigates the NLOS bias problem by estimating the NLOS path length directly and is only limited in performance by the system parameters like noise figure and Bandwidth.

### v2

@@ IV-A ToF based range measurements of the first arriving path
Here, $\gamma_{k}(\theta_{tx,k},\theta_{rx,k})$ is the complex gain for the received signal due to the $k^{th}$ MPC and includes the effects of the transmit and receive side antennas, transmitted signal power and also the path gain $h_{k}$. For further analysis, the receiver noise $n(t)$ is assumed to be Gaussian. Of the $K$ MPCs in the received signal, let the indices $k=0$ and $k=1$ correspond to the two diffraction MPCs. Therefore, the remaining $K$-$2$ MPCs consist of those MPCs that contain more than one interaction with the environment. If the electric field for the transmit signal is parallel to the horizontal diffracting window edges, Theorem 1 applies. Therefore, it leads to the upper edge diffraction MPC dominating over the lower edge diffraction MPC. In a system this could be achieved by mounting the linear polarized antennas horizontally, instead of the conventional vertically mounted implementation. Thus, the received signal can be written as
$\begin{split}&r(t)\approx\underbrace{\gamma_{0}(\theta_{tx,0},\theta_{rx,0})p(t-\tau_{0})}_{\text{dominating diffraction MPC}}\\ &\hskip 50.00008pt+\underbrace{\sum_{k=2}^{K-1}\gamma_{k}(\theta_{tx,k},\theta_{rx,k})p(t-\tau_{k})}_{\text{MPC-4}}+n(t)\\ \end{split}$ (17)
Next, we assume resolvability between the MPCs i.e., $|\tau_{i}-\tau_{j}|>T_{p},\;\;\forall\;\;i\neq j$. It can be shown with this assumption, that the CRLB on the RMSE for estimating the propagation delay of the first arriving path $\tau_{0}$ is given by [41, 42]
$\begin{split}\text{CRLB}(\tau_{0})=\frac{1}{\sqrt{8\pi^{2}\beta^{2}\text{SNR}}}=\sqrt{\frac{N_{0}}{8\pi^{2}\beta^{2}(\gamma_{0}(\theta_{rx,0},\theta_{tx,0}))^{2}}}.\end{split}$ (18)
Here, $\gamma_{0}(\theta_{rx,0},\theta_{tx,0})$ is the amplitude corresponding to the first arriving path and includes the transmit and receive side antenna gains, SNR is receive side signal-to-noise ratio (linear), and $\beta$ is the bandwidth of the transmitted signal (Hz). Since this first arriving path corresponds to the dominating diffraction path, we have $\tau_{0}=\frac{p}{c}$. Here, $\tau_{0}$ is the propagation delay of the first arriving path, and the path length of the dominating diffraction path $p$ is given by eq. (12) and $c$ is the speed of light. Drawing insights from the raytracing analysis in Appendix -D and eq. (18), we can improve the estimate of $p(\bm{\alpha})$ by maximizing the receiver side antenna gain $\alpha_{rx,k}(\theta_{rx,k})$ for AoA $\theta_{rx,k}=90$°. In other words, we could improve the ranging performance if we orient the receiver side antenna such that the antenna gain is maximum towards the windows. Further improvement to the ranging estimate can be made by scanning the transmit side beam to maximize the receiver side power of the dominant diffracting MPC.
@@ IV-B Fisher Information analysis of the 3D position estimate
In this section, we examine the conditions under which we can perform 3D NLoS localization using the ToF-based ranging estimates corresponding to the dominating diffraction MPC using Fisher Information analysis and then calculate the lower bound to the 3D position error. In Fig. 4, we have ‘$M$’ UAV anchors in front of the building. The node is located on an arbitary building floor which has two diffracting edges separated by a vertical distance $w$. We assume that the origin is located in the middle bottom of the front wall of the building at the intersection of the front wall with the ground, the x axis oriented along the diffracting edges and the z-axis pointed along the vertical direction as in Fig.,4.
@@
Each anchor indexed by ‘$j$’ is placed at known coordinates $\bm{X}^{a}_{j}=[x_{a,j},y_{a,j},z_{a,j}]^{T}$ relative to the origin, while the node is located on an arbitrary floor of the building at $\bm{\alpha}=[x_{n},y_{n},z_{n}]^{T}$. Assuming that the anchors transmit x-polarized signals, each ToF-based range measurement corresponds to the path length of the respective dominant diffraction MPC between each anchor $j$ and the node. With a slight change in notation, we denote the path length of the respective dominating diffraction MPC as $p_{j}(\bm{\alpha})$ obtained using eq. (12), with the subscript ‘$j$’ representing the $j^{\text{th}}$ anchor. Consequently, we obtain $M$ noisy range measurements from which we need to estimate the 3D position of the node. Assuming the 3D position of the node $\bm{\alpha}$ as the vector of the three parameters to be estimated, its corresponding Fisher Information Matrix (FIM) $\bm{J}_{\bm{\alpha}}\in\mathbb{R}^{3\times 3}$ [43], [44] and can be expressed as
$\bm{J}_{\bm{\alpha}}=c^{2}\bm{J}^{T}\bm{J}_{\tau_{0}}\bm{J}.$ (19)
Since we assume the noise in our range measurements is independent across the $M$ anchors, $\bm{J}_{\tau_{0}}=diag(\lambda_{0},\cdots,\lambda_{M-1})$ is a diagonal matrix where $\lambda_{j}=\frac{1}{8\pi^{2}\beta^{2}SNR_{j}}$ represents the inverse of the CRLB of the path delay estimate of the first arriving path between the $j^{th}$ anchor and the node. The Jacobian $\bm{J}\in\mathbb{R}^{M\times 3}$ is the transform from the ranging measurement space $\in\mathbb{R}^{M}$ to the unknown parameter to be estimated space $\in\mathbb{R}^{3}$ and is given by
$\bm{J}=\begin{bmatrix}\frac{\partial{p_{0}(\bm{\alpha})}}{\partial x_{n}}&\frac{\partial{p_{0}(\bm{\alpha})}}{\partial y_{n}}&\frac{\partial{p_{0}(\bm{\alpha})}}{\partial z_{n}}\\ \vdots&\vdots&\vdots\\ \frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial x_{n}}&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial y_{n}}&\frac{\partial{p_{M-1}(\bm{\alpha})}}{\partial z_{n}}\\ \end{bmatrix}.$ (20)
@@
The FIM $\bm{J}_{\bm{\alpha}}\in\mathbb{R}^{3\times 3}$ defined in eq. (19) is invertible if and only if $rank(\bm{J})=3$. Here, the Jacobian Matrix $\bm{J}\in\mathbb{R}^{M\times 3}$ is defined in eq. (20) from the ToF ranging measurement space $\mathbb{R}^{M}$ to the 3D coordinate space $\bm{\alpha}\in\mathbb{R}^{3}$.
@@
$\sqrt{\bm{A}}$ is the elementwise square root operation of a matrix $\bm{A}$. We also know that multiplying a matrix with a non-zero diagonal matrix does not change its rank. Furthermore, by using $rank(\bm{A}\bm{A}^{T})=rank(\bm{A})$ [46], we obtain $rank(\bm{J}_{\bm{\alpha}})=rank(\bm{J})$. Now, $\bm{J}_{\bm{\alpha}}$ is invertible if and only if $rank(\bm{J}_{\bm{\alpha}})=3$. Hence, we obtain the desired result. ∎
@@
  • It is necessary but not sufficient to have at least four anchors. This is because, for $rank(\bm{J})=3$, it is necessary but not sufficient to have $M\geq 3$.
@@
This result demonstrates that we no longer require detailed position knowledge about the diffraction edges which is an improvement over the conference version of this paper [1] that relied on the path model in Lemma 1. Essentially by incorporating the fact that the diffraction occurs on the same floor as the UE and that the vertical coordinate of the diffraction point is very close to that of the node, we are able to reduce one degree of freedom from the path model in Lemma 1 to Corollary 1.
To illustrate that having $M=3$ anchors is not a sufficient condition for 3D positioning, consider a scenario where we have three anchors such that their x-coordinate $x_{a,j}=0\;\;\forall\;\;j\in[0,2]$. Observe that as the node position is varied inside the building such that it approaches the plane $X=0$, i.e., $x_{n}\to 0$, for each anchor, we have the x-coordinate of the corresponding diffraction point $q_{x}\to 0$. Notice in the first row of the Jacobian $\bm{J}$, defined in eq. (-E), the term $\frac{\partial p_{j}{(\bm{\alpha}})}{\partial x_{n}}\to 0\;\;\forall j\in[0,2]$. Therefore, for $x_{n}=0$, $rank(\bm{J})\neq 3$, and by using Theorem 2, we cannot estimate the 3D position of the node for this scenario.
The diffracting edges are located parallel to the X-axis at $Y=0$. Observe the x-coordinates of the end points of the diffracting edges correspond to the width of the building which is required to be known a priori whereas the height of the diffracting edges is not required to be known. This is because the node and the edge from which diffraction takes place are located on the same building floor. Therefore, by estimating the node location leads to an estimate of the height of the edge since the diffracting edge is located at a distance $0.5w$ above the node. Note Corollary 3 leads to the following observation - we need position knowledge of the anchors in local building coordinates. These local coordinates are defined with respect to the origin located on the diffracting edges as shown in Fig. 4.
@@
V Simulation Results
In this section, we first describe a simulation strategy to generate noisy range estimates from raytracing simulation data. Using these range measurements, we then propose a non-linear least squares algorithm called ‘D-NLS’ for 3D positioning based on our diffraction path model. For comparison, we implement two alternative approaches: (a) Linear Least Squares (LLS), which assumes that all range measurements are LoS, and (b) the Iterative Parallel Projection Algorithm (IPPA) for positioning in NLoS scenarios [7, 8]. The IPPA is a set-theoretic method that mitigates the impact of NLoS bias by utilizing information about whether a range measurement is LoS or NLoS. Los/NLos identification can be achieved by using techniques in [10].
From our raytracing simulation in Appendix -D, using a thresholding logic we pick the first arriving path between each candidate Rx location and each UAV anchor. Observe from Table VI the first arriving path across receivers are majorly the diffraction paths i.e. MPCs present in MPC-3. These should have a true path length derived in Corollary 1. Our measurement model for the range measurements at each Rx location $\bm{\alpha}$ is
$\bm{r}=\mathbf{p(\bm{\alpha})+\mathbf{n}}.$ (21)
Here, $\bm{r}=[r_{0},\cdots,r_{j},\cdots,r_{M-1}]^{T}\in\mathbb{R}^{M}$ is the vector of the noisy range measurements from $M=4$ anchors, $\mathbf{p(\bm{\alpha})}=[p_{0}(\bm{\alpha}),\cdots,p_{j}(\bm{\alpha}),\cdots,p_{M-1}(\bm{\alpha})]^{T}$, where $p_{j}(\bm{\alpha})$ is the diffraction path length between the $j^{\text{th}}$ anchor and node obtained using Corollary 1. Now the uncertainty in range measurements is limited by both SNR and signal Bandwidth. Hence from Section IV-A, the noise vector is simulated as $\bm{n}=[n_{0},\cdots,n_{M-1}]$, where $n_{j}\sim\mathcal{N}(0,\sigma_{j}^{2})$, $\sigma_{j}=\frac{c}{\sqrt{8\pi^{2}\beta^{2}\text{SNR}_{j}}}$ is the uncertainty in the range measurements for $j^{\text{th}}$ anchor.
@@ V Simulation Results
Anchor $1$ pos. $\bm{X}_{0}^{a}=[-15m,-20m,1m]^{T}$
Anchor $2$ pos. $\bm{X}_{1}^{a}=[0m,15m,0m]^{T}$
Anchor $3$ pos. $\bm{X}_{2}^{a}=[0m,-7m,4m]^{T}$
Anchor $4$ pos. $\bm{X}_{3}^{a}=[15m,-20m,1m]^{T}$
Signal Bandwidth $\beta=400$ MHz
Noise Floor $N_{0}$=-87 dBm
Rx processing gain 20dB
V-A 3D positioning algorithm based on the diffraction path model
From the measurement model in eq. (21), we have the noisy range measurements $\bm{r}$ and the UAV anchor positions from which we need to obtain a position estimate $\bm{\hat{\alpha}}$. This can be achieved by minimizing the least squared error between the range measurements and the diffraction path model as
$\bm{\hat{\alpha}}=\underset{\bm{\alpha}}{\mathrm{\text{arg min}}}\sum_{j=1}^{M}|r_{j}-{p_{j}}(\bm{\alpha})|^{2}.$
Since the diffraction path length model $p_{j}(\bm{\alpha})$ is a non-linear function we utilize a Gauss-Newton approach [43] to derive an iterative estimator referred to as D-NLS (Case-1). Using subscript $m$ to denote the iteration index the iterative update steps are
$\bm{\hat{\alpha}}_{m+1}=\bm{\hat{\alpha}}_{m}+(\bm{J}_{m}^{T}\bm{J}_{m})^{-1}\bm{J}_{m}^{T}(\bm{r}-\bm{p}(\bm{\hat{\alpha}}_{m})).$ (22)
Here $\bm{J}_{m}$ is the Jacobian matrix defined in eq. (20) for the $m^{\text{th}}$ iteration. In the analysis until now, we had assumed that the node is located at the midpoint of the building floor as proposed in Assumption 1. However, this may not necessarily hold true as the indoor user might lay prone on the floor or the receiver may be mounted at a different height. We account for this by assuming that there is a distance offset $\Delta$ between the node’s z-coordinate and the upper diffraction edge, i.e.
$z_{n}=z_{e}^{u}-0.5w+\Delta.$ (23)
Clearly, $\Delta$ is upper bounded by half the height of the building floor, and by including this offset in our path model in our least squares formulation, we obtain positioning performance which includes a slight model mismatch. Note, $\Delta$ is not estimated as part of the positioning algorithm. It merely represents the amount of model mismatch between our assumption that the z-coordinate of the node is at the mid-point of the building floor and reality. Thus we compare performance with and without Assumption 1 by comparing D-NLS (Case-1) vs D-NLS (Case-2). Note, in D-NLS (Case-2) which includes a model mismatch sets an upper bound to our proposed diffraction-based positioning technique.
V-B Results and discussion
To obtain the theoretical lower bound to the performance, we use the Position Error Bound (PEB) obtained as $\text{PEB}=\sqrt{\text{Tr}({\bm{J_{\alpha}}}^{-1})}$ where $\text{Tr}{(\bm{A})}$ is the trace of matrix $\bm{A}$ and $\bm{J_{\alpha}}$ is the FIM from Section IV-B. Similarly, the bound on the z-axis is defined as $\sqrt{{(\bm{J}^{-1}_{\bm{\alpha}})}^{[3,3]}}$, where the notation $\bm{A}^{[i,j]}$ represents the element in the $i^{th}$ row and $j^{th}$ column of the matrix $\bm{A}$. Since we have the ground-truth positions of Rx from the raytracing study, we can calculate the 3D RMSE=$\sqrt{(\bm{\alpha}-\bm{\hat{\alpha}})^{T}(\bm{\alpha}-\bm{\hat{\alpha}})}$ for each positioning approach. Similarly, we can calculate the RMSE in z-axis estimate. Finally, we present the CDF of the RMSE and PEB in Fig. 5. The LLS approach serves as the baseline comparison but performs poorly because it assumes LoS conditions, and due to poor anchor geometry since they are located on one side of the building. IPPA is able to improve the positioning performance but due to its reliance on modeling the path length using Euclidean distance, it is limited in its effectiveness at mitigating the positioning errors caused by the NLoS bias in the range measurements. D-NLS (Case-1), our proposed NLS technique based on the diffraction path model including Assumption 1, offers the best algorithmic performance, but does not meet the CRLB. This is because of two main reasons (a) model mismatch (b) the NLS approach is sensitive to initialization values and sometimes converges to local minima [44]. The first arriving paths are not always diffraction paths (i.e., not included in MPC-3, as defined in Appendix -D). This may occur either because these paths do not exist at certain Rx locations or because their SNR is extremely low, resulting in missed detections by our first-arriving path estimator. Therefore, in these cases, we assume that the range measurements are due to diffraction in eq. (21) would be incorrect and this would be a case of mismatch of the model. Note that in our simulation, we also simulate relaxing Assumption 1 in D-NLS(Case-2) , which assumes that the node positions are at the midpoint of the building floor. To account for this, we introduced an offset $\Delta=1.5m$, corresponding to half the height of a building floor, in our diffraction path model according to eq. (23). Although including this offset slightly degrades positioning performance as shown in D-NLS (Case-2), our proposed technique still achieves better positioning accuracy in NLoS scenarios compared to previous methods.
(a) 3D positioning Performance.
(b) Z-axis positioning Performance.
Figure 5: We evaluate the positioning performance in NLoS conditions by comparing the lower bound derived using the CRLB for our proposed diffraction-inspired positioning technique (D-NLS), both with and without Assumption 1 (Case-1 vs. Case-2). Our approach demonstrates improved positioning performance compared to both the prior art in NLoS positioning (IPPA) and the baseline method (LLS) for (a) 3D positioning and (b) z-axis positioning accuracy.